| | | AGuidetoPreparing Documentswith−ms M.E.Lesk BellLaboratoriesAugust1978 ThisguidegivessomesimpleexamplesofdocumentpreparationonBellLabscomputers, emphasizingtheuseofthe−−mmssmacropackage. Itenormouslyabbreviatesinformationin 1.TTyyppiinnggDDooccuummeennttssoonnUUNNIIXXaannddGGCCOOSS,,by M.E.Lesk; 2.TTyyppeesseettttiinnggMMaatthheemmaattiiccss−−UUsseerr’’ssGGuuiiddee,, byB.W.KernighanandL.L.Cherry;and 3.TTbbll−−AAPPrrooggrraammttooFFoorrmmaattTTaabblleess,,byM. E.Lesk. ThesememosareallincludedintheUUNNIIXXPPrroo-- ggrraammmmeerr’’ssMMaannuuaall,,VVoolluummee22..Thenewuser shouldalsohaveAATTuuttoorriiaallIInnttrroodduuccttiioonnttootthhee UUNNIIXXTTeexxttEEddiittoorr,,byB.W.Kernighan. Formoredetailedinformation,readAAddvvaanncceedd EEddiittiinnggoonnUUNNIIXXandAATTrrooffffTTuuttoorriiaall,,byB.W. Kernighan,and(forexperts)NNrrooffff//TTrrooffffRReeffeerr-- eenncceeMMaannuuaallbyJ.F.Ossanna.Informationon relatedcommandsisfound(forUNIXusers)in UUNNIIXXffoorrBBeeggiinnnneerrssbyB.W.Kernighanand theUUNNIIXXPPrrooggrraammmmeerr’’ssMMaannuuaallbyK.ThompsonandD.M.Ritchie. Contents ATM....................2 Areleasedpaper.............3 Aninternalmemo,andheadings...4 Lists,displays,andfootnotes......5 Indents,keeps,anddoublecolumn..6 Equationsandregisters.........7 Tablesandusage.............8 Throughouttheexamples,inputisshownin thisHelveticasansseriffont whiletheresultingoutputisshownin thisTimesRomanfont. UNIXDocumentno.1111 2 CCoommmmaannddssffoorraaTTMM .TM1978-5b39999999999-11 .NDApril1,1976 .TL TheRoleoftheAllenWrenchinModern Electronics .AU"MH2G-111"2345 J.Q.Pencilpusher .AU"MH1K-222"5432 X.Y.Hardwired .AI .MH .OK Tools Design .AB Thisabstractshouldbeshortenoughto fitonasinglepagecoversheet. Itmustattractthereaderintosendingfor thecompletememorandum. .AE .CS10212567 .NH Introduction. .PP Nowthefirstparagraphofactualtext... ... Lastlineoftext. .SGMH-1234-JQP/XYH-unix .NH References... Commandsnotneededinaparticularformatare ignored. BellLaboratoriesBellLaboratoriesCoverSheetforTM ThisinformationisforemployeesofBellLaboratories.(GEI13.9-3) Title-TThheeRRoolleeoofftthheeAAlllleennWWrreenncchh iinnMMooddeerrnnEElleeccttrroonniiccss Date-AApprriill11,,11997766 TM-11997788--55bb33 OtherKeywords-TToooollss DDeessiiggnn AuthorLocationExt.ChargingCase-9999999999 JJ..QQ..PPeenncciillppuusshheerrMMHH22GG--11111122334455 FilingCase-99999a99999a XX..YY..HHaarrddwwiirreeddMMHH11KK--22222255443322 ABSTRACT Thisabstractshouldbeshortenoughto fitonasinglepagecoversheet.Itmust attractthereaderintosendingforthecompletememorandum. PagesText10Other2Total12 No.Figures5No.Tables6No.Refs.7||||| E-1932-U(6-73)SEEREVERSESIDEFORDISTRIBUTIONLIST | | | | | | 3 AARReelleeaasseeddPPaappeerrwwiitthhMMaatthheemmaattiiccss .EQ delim$$ .EN .RP ...(asforaTM) .CS10212567 .NH Introduction .PP Thesolutiontothetorquehandleequation .EQ(1) sumfrom0toinfF(xsubi)=G(x) .EN isfoundwiththetransformation$x=rhoover theta$where$rho=Gprime(x)$and$theta$ isderived... TThheeRRoolleeoofftthheeAAlllleennWWrreenncchh iinnMMooddeerrnnEElleeccttrroonniiccss J.Q.Pencilpusher X.Y.Hardwired BellLaboratories MurrayHill,NewJersey07974 ABSTRACT Thisabstractshouldbeshortenoughtofitona singlepagecoversheet.Itmustattractthereader intosendingforthecompletememorandum. April1,1976 TThheeRRoolleeoofftthheeAAlllleennWWrreenncchh iinnMMooddeerrnnEElleeccttrroonniiccss J.Q.Pencilpusher X.Y.Hardwired BellLaboratories MurrayHill,NewJersey07974 11..IInnttrroodduuccttiioonn Thesolutiontothetorquehandleequation ∞0ΣF(xi)=G(x)(1) isfoundwiththetransformationx=ρθwhereρ=G′(x)and θisderivedfromwell-knownprinciples.

4 AAnnIInntteerrnnaallMMeemmoorraanndduumm .IM .NDJanuary24,1956 .TL The1956ConsentDecree .AU Able,Baker& Charley,Attys. .PP Plaintiff,UnitedStatesofAmerica,havingfiled itscomplainthereinonJanuary14,1949;the defendantshavingappearedandfiledtheir answertosuchcomplaintdenyingthe substantiveallegationsthereof;andtheparties, bytheirattorneys,... BBeellllLLaabboorraattoorriieess Subject:

TThhee11995566CCoonnsseennttDDeeccrreeee date:

JJaannuuaarryy2244,,11995566 from:

AAbbllee,,BBaakkeerr&& CChhaarrlleeyy,,AAttttyyss.. Plaintiff,UnitedStatesofAmerica,havingfileditscomplainthereinonJanuary14,1949;thedefendantshaving appearedandfiledtheiranswertosuchcomplaintdenying thesubstantiveallegationsthereof;andtheparties,bytheir attorneys,havingseverallyconsentedtotheentryofthis FinalJudgmentwithouttrialoradjudicationofanyissuesof factorlawhereinandwithoutthisFinalJudgmentconstitutinganyevidenceoradmissionbyanypartyinrespectofany suchissues; Now,thereforebeforeanytestimonyhasbeentakenherein, andwithouttrialoradjudicationofanyissueoffactorlaw herein,andupontheconsentofallpartieshereto,itis hereby Ordered,adjudgedanddecreedasfollows: II..[[SShheerrmmaannAAcctt]] ThisCourthasjurisdictionofthesubjectmatterherein andofallthepartieshereto.Thecomplaintstatesaclaim uponwhichreliefmaybegrantedagainsteachofthedefendantsunderSections1,2and3oftheActofCongressof July2,1890,entitled‘‘Anacttoprotecttradeandcommerce againstunlawfulrestraintsandmonopolies,’’commonly knownastheShermanAct,asamended. IIII..[[DDeefifinniittiioonnss]] ForthepurposesofthisFinalJudgment: (a)‘‘Western’’shallmeanthedefendantWesternElectric Company,Incorporated. Otherformatspossible(specifybefore.TL)are:.MR (‘‘memoforrecord’’),.MF(‘‘memoforfile’’),.EG (‘‘engineer’snotes’’)and.TR(ComputingScience Tech.Report). HHeeaaddiinnggss .NH .SH Introduction.AppendixI .PP .PP texttexttexttexttexttext 11..IInnttrroodduuccttiioonnAAppppeennddiixxII texttexttexttexttexttext | | | 5 AASSiimmpplleeLLiisstt .IP1. J.PencilpusherandX.Hardwired, .I ANewKindofSetScrew, .R Proc.IEEE .B75 (1976),23-255. .IP2. H.NailsandR.Irons, .I FastenersforPrintedCircuitBoards, .R Proc.ASME .B23 (1974),23-24. .LP(terminateslist) 1.J.PencilpusherandX.Hardwired,ANewKind ofSetScrew,Proc.IEEE7575(1976),23-255. 2.H.NailsandR.Irons,FastenersforPrintedCircuitBoards,Proc.ASME2323(1974),23-24. DDiissppllaayyss texttexttexttexttexttext .DS andnow forsomething completelydifferent .DE texttexttexttexttexttext hobokenharrisonnewarkrosevilleavenuegrovestreet eastorangebrickchurchorangehighlandavenue mountainstationsouthorangemaplewoodmillburn shorthillssummitnewprovidence andnow forsomething completelydifferent murrayhillberkeleyheightsgillettestirlingmillington lyonsbaskingridgebernardsvillefarhillspeapack gladstone Options:.DSL:left-adjust;.DSC:line-by-linecenter; .DSB:makeblock,thencenter. FFoooottnnootteess Amongthemostimportantoccupants oftheworkbencharethelong-nosedpliers. Withoutthesebasictools* .FS *AsfirstshownbyTiger&Leopard (1975). .FE | | | fewassembliescouldbecompleted.Theymay lackthepopularappealofthesledgehammer Amongthemostimportantoccupantsoftheworkbencharethelong-nosedpliers.Withoutthesebasic tools*fewassembliescouldbecompleted.Theymay lackthepopularappealofthesledgehammer *AsfirstshownbyTiger&Leopard(1975).

6 MMuullttiipplleeIInnddeennttss Thisisordinarytexttopointout themarginsofthepage. .IP1. Firstlevelitem .RS .IPa) Secondlevel. .IPb) Continuedherewithanothersecond levelitem,butsomewhatlonger. .RE .IP2. Returntopreviousvalueofthe indentingatthispoint. .IP3. Another line. Thisisordinarytexttopointoutthemarginsofthe page. 1.Firstlevelitem a)Secondlevel. b)Continuedherewithanothersecondlevel item,butsomewhatlonger. 2.Returntopreviousvalueoftheindentingatthis point. 3.Anotherline. KKeeeeppss Linesbracketedbythefollowingcommandsarekept together,andwillappearentirelyononepage: .KSnotmoved.KFmayfloat .KEthroughtext.KEintext DDoouubblleeCCoolluummnn .TL TheDeclarationofIndependence .2C .PP Wheninthecourseofhumanevents,itbecomes necessaryforonepeopletodissolvethepoliticalbonds whichhaveconnectedthemwithanother,andtoassume amongthepowersoftheearththeseparateandequal stationtowhichthelawsofNatureandofNature’sGod entitlethem,adecentrespecttotheopinionsof TheDeclarationofIndependence Wheninthecourseof humanevents,itbecomes necessaryforonepeople todissolvethepolitical bondswhichhaveconnectedthemwithanother,andtoassumeamong thepowersoftheearth theseparateandequal stationtowhichthelaws ofNatureandofNature’s | | | Godentitlethem,adecentrespecttotheopinionsofmankindrequires thattheyshoulddeclare thecauseswhichimpel themtotheseparation. Weholdthesetruths tobeself-evident,thatall menarecreatedequal, thattheyareendowedby theircreatorwithcertain unalienablerights,that amongthesearelife,liberty,andthepursuitof happiness.Thattosecuretheserights,governmentsareinstituted amongmen, | | | 7 EEqquuaattiioonnss Adisplayedequationismarked withanequationnumberattherightmargin byaddinganargumenttotheEQline: .EQ(1.3) xsup2overasup2˜ =˜ sqrt{pzsup2+qz+r} .EN Adisplayedequationismarkedwithanequation numberattherightmarginbyaddinganargumentto theEQline: x2a2=√pz2+qz+r(1.3) .EQI(2.2a) boldVbarsubnu˜ =˜ left[pile{aabovebabove c}right]+left[matrix{col{A(11)above. above.}col{.above.above.}col{.above. aboveA(33)}}right]cdotleft[pile{alpha abovebetaabovegamma}right] .EN VVν=[abc]+ A(11).......A(33) . αβγ (2.2a) .EQL Fhat(chi)˜ mark=˜ |delV|sup2 .EN .EQL lineup=˜ {left({partialV}over{partialx}right)} sup2+{left({partialV}over{partialy}right)}sup 2˜ ˜ ˜ ˜ ˜ ˜ lambda->inf .EN ˆ F(χ)=|∇V|2 =(∂V∂x)2+(∂V∂y)2λ→∞ $adot$,$bdotdot$,$xitildetimesyvec$: .a,..b,˜ ξ×→y.(withdelim$$on,seepanel3). Seealsotheequationsinthesecondtable,panel8. SSoommeeRReeggiisstteerrssYYoouuCCaannCChhaannggee Linelength .nrLL7i Titlelength .nrLT7i Pointsize .nrPS9 Verticalspacing .nrVS11 | | | Columnwidth .nrCW3i Intercolumnspacing .nrGW.5i Margins−headandfoot .nrHM.75i .nrFM.75i Paragraphindent .nrPI2n Paragraphspacing .nrPD0 Pageoffset .nrPO0.5i Pageheading .dsCHAppendix (center) .dsRH7-25-76 (right) .dsLHPrivate (left) Pagefooter .dsCFDraft .dsLF .dsRFsimilar Pagenumbers .nr%3 8 TTaabblleess .TS(T◯indicatesatab) allbox; css ccc nnn. AT&TCommonStock YearT◯PriceT◯Dividend 1971T◯41-54T◯$2.60 2T◯41-54T◯2.70 3T◯46-55T◯2.87 4T◯40-53T◯3.24 5T◯45-52T◯3.40 6T◯51-59T◯.95* .TE *(firstquarteronly)

AT&TCommonStock YearPriceDividend 197141-54$2.60 241-542.70 346-552.87 440-533.24 545-523.40 651-59.95* *(firstquarteronly) Themeaningsofthekey-lettersdescribingthealign-mentofeachentryare: ccenternnumerical rright-adjustasubcolumn lleft-adjustsspanned Theglobaltableoptionsarecenter,expand,box, doublebox,allbox,tab(xx)andlinesize(nn). .TS(withdelim$$on,seepanel3) doublebox,center; cc ll. NameT◯Definition .sp GammaT◯$GAMMA(z)=intsub0supinf\ tsup{z-1}esup-tdt$ SineT◯$sin(x)=1over2i(esupix-esup-ix)$ ErrorT◯$romanerf(z)=2oversqrtpi\ intsub0supzesup{-tsup2}dt$ BesselT◯$Jsub0(z)=1overpi\ intsub0suppicos(zsintheta)dtheta$ ZetaT◯$zeta(s)=\ sumfromk=1toinfksup-s˜ ˜ (Re˜ s>1)$ .TE NameDefinition GammaΓ(z)=∫∞0tz−1e−tdt Sinesin(x)=12i(eix−e−ix) Errorerf(z)=2√π∫z0e−t2dt BesselJ0(z)=1π∫π0cos(zsinθ)dθ Zetaζ(s)=∞k=1Σk−s(Res>1) UUssaaggee | | | Documentswithjusttext: troff-msfiles Withequationsonly: eqnfiles|troff-ms Withtablesonly: tblfiles|troff-ms Withbothtablesandequations: tblfiles|eqn|troff-ms TheabovegeneratesSTAREoutputonGCOS:replace −stwith−phfortypesetteroutput. | | | 70504 | | |